OpenStudy (anonymous):

evaluate cos^-1(sqrt(2)sinπ/4)

zepdrix (zepdrix):

\[\large \arccos\left[\sqrt2 \color{royalblue}{\sin\left(\frac{\pi}{4}\right)}\right]\] So let's work from the inside outward. Start with the blue part. Do you remember what this special angle produces?

OpenStudy (anonymous):

No clue :/

zepdrix (zepdrix):

Booooooo!

zepdrix (zepdrix):

If you're working on problems like this, you're suppose to know that by now :c Try to take some time and commit the special angles and measurements on the unit circle to memory.

OpenStudy (anonymous):

So would the answer for the blue part be a angle in degrees?

OpenStudy (anonymous):

Oh never mind! is it square root(2)/2?

zepdrix (zepdrix):

No, we're taking the sine OF an angle. It will produce a length. Yes, good. That's the one :)

OpenStudy (anonymous):

ok so the blue part is square root (2)/2 whats next?

zepdrix (zepdrix):

\[\large \arccos\left[\sqrt2 \color{royalblue}{\frac{\sqrt2}{2}}\right]\]

zepdrix (zepdrix):

Hmm it looks like that inside part will simplify a bit, doesn't it? :o

OpenStudy (anonymous):

2/2 which equals 1?

zepdrix (zepdrix):

\[\large \arccos\left[1\right]\]

zepdrix (zepdrix):

Cool looks like we're almost done! :)

OpenStudy (anonymous):

Ok so would you just type that into the calculator to get the answer?

zepdrix (zepdrix):

For this one, it turns out yes you can. It will give you a nice easy answer. But normally that might not work. Here is how you should think of it.\[\large \arccos\left[1\right]=\theta\]We can rewrite this as the cosine, like so,\[\large \cos \theta=1\] What special angle produces 1 in the cosine? :O

OpenStudy (anonymous):

360 degrees or 2π?

zepdrix (zepdrix):

Yes, but there is a better value we can use :) What is 2pi equivalent to? Spin backward one full spin around the unit circle, where do you land? :O

OpenStudy (anonymous):

just π?

zepdrix (zepdrix):

No silly! You only spun around half way! :3

OpenStudy (anonymous):

Sorry! 0?

zepdrix (zepdrix):

Yah sounds good c: \(\large \cos \theta=1 \qquad \qquad \rightarrow \qquad \qquad \theta=0\) This is actually a bit important. Because the INVERSE cosine, can't produce an angle of 2pi. It has to produce an angle between 0 and pi. So it's important that you chose the 0, not the 2pi angle.

OpenStudy (anonymous):

zepdrix i love your photo

zepdrix (zepdrix):

haha thanks XD Yours is a little scary <:O hehe

OpenStudy (anonymous):

Ok thank you so much again! i have a feeling im gonna be thanking you a lot today haha

zepdrix (zepdrix):

:3